# BlasBench: An Open Benchmark for Irish Speech Recognition github.com/jyoutir/blasbench

### 12 Apr 2026 v1

@@ Abstract
No open Irish-specific benchmark compares end-user ASR systems under a shared Irish-aware evaluation protocol. To solve this, we release BlasBench, an open evaluation harness with Irish-aware text normalisation that preserves fadas, lenition, and eclipsis. We benchmark 12 systems across four architecture families on Common Voice ga-IE and FLEURS ga-IE. All Whisper variants exceed 100% WER. The best open model (omniASR LLM 7B) achieves 30.65% WER on Common Voice and 39.09% on FLEURS. We noticed models fine-tuned on Common Voice lose 33–43 WER points on FLEURS, revealing a generalisation gap that is invisible to single-dataset evaluation.
@@
Jyoutir Rajthanks: Corresponding author. Independent Researcher jyoutirraj@gmail.com          John Conway threefold.eco johnconway114@gmail.com
@@ 1 Introduction
Irish speech recognition does not work well. The commercial system we test, Microsoft Azure, gets 22.2% WER on read speech (Common Voice ga-IE) and 57.5% WER on the more diverse FLEURS ga-IE and every Whisper variant we evaluate (Radford et al., 2023) produces more tokens than the reference contains—over 100% WER on both datasets (Tables 12).
The problem is not that nobody has worked on Irish ASR. The ABAIR group at Trinity College Dublin has produced a deep body of work culminating in their Fotheidil system at 10.9% WER (Lonergan et al., 2025). Qian et al. (2024) fine-tune Whisper on Irish. Community wav2vec2 fine-tunes exist on HuggingFace (Faste, 2022). The IWSLT shared task has produced Irish–English speech translation systems (Ojha and McCrae, 2024; Moslem, 2024). Multilingual benchmarks—FLEURS (Conneau et al., 2022a), MMS (Pratap et al., 2023), ML-SUPERB (Shi et al., 2023)—include Irish among dozens or hundreds of languages.
The problem is that these efforts cannot be compared or reliably reproduced. ABAIR’s best result is on private MíleGlór data, and while they also report on Common Voice and FLEURS-R, they benchmark only their own systems. Community models self-report on different Common Voice versions with different normalisers. Multilingual benchmarks include Irish without Irish-specific normalisation or analysis. [^1] ML-SUPERB and XTREME-S (Conneau et al., 2022b) include Irish data points but do not compare end-user ASR systems on Irish specifically. We did not identify a prior open, Irish-specific comparison of end-user ASR systems across architecture families under a shared Irish-aware evaluation protocol.
We present BlasBench, [^2] Harness: github.com/jyoutir/blasbench. Per-utterance predictions and aggregate results for every run in this paper are released as v0.1.0. an open benchmark and evaluation harness for Irish ASR. The benchmark is the primary contribution; the 12-model comparison is its first use case. Reproducible, language-aware scoring is also a prerequisite for credible ASR comparison and for any automated research pipeline (Karpathy, 2026) that needs a stable signal. It matters especially for Irish, whose orthography encodes grammar and phonemic contrasts that language-agnostic normalisers erase. Dedicated low-resource benchmarks for Urdu (Akhtar et al., 2024) and Scottish Gaelic (Klejch et al., 2025) make the same case; BlasBench brings it to Irish.
@@ 2.3 Evaluation pipeline
    Scoring. WER and CER are computed via global aggregation: total substitutions, insertions, and deletions across all utterances divided by total reference units. This avoids the bias of per-sentence averaging (Akhtar et al., 2024).
@@
    Confidence intervals. Bootstrap 95% CIs (Bisani and Ney, 2004) are computed with 1,000 resamples at a fixed seed (42), resampling utterance-level error counts and recomputing the global aggregate.
@@ 2.4 Irish-aware normalisation
Standard ASR normalisers lowercase text, strip punctuation, and collapse whitespace. Recent work on multilingual ASR evaluation shows that language-agnostic normalisation can materially distort reported error rates (Manohar and Pillai, 2024). For Irish, it destroys linguistically significant information.
@@
ABAIR’s normalisation module handles digit and acronym expansion (Lonergan et al., 2025) but is not released standalone. To our knowledge, BlasBench ships the first standalone open-source Irish ASR evaluation normaliser released specifically for Irish ASR scoring; unreleased internal tools may exist.
@@ 3 Systems Evaluated
Whisper (Radford et al., 2023): Encoder-decoder transformer. Irish is absent from Whisper’s language list up to large-v2 and appears in the large-v3 tokenizer with no reported Irish training audio,333ga is absent from the LANGUAGES dictionary in whisper/tokenizer.py for v1/v2 and carries no training-data disclosure for v3. so results here are effectively zero-shot. We test medium (769M), large-v2 (1.5B), large-v3 (1.5B), and large-v3-turbo (809M).
@@
Meta multilingual: MMS-1B-All (Pratap et al., 2023) (1,100+ languages); omniASR LLM 300M and 7B, wav2vec2 encoders with LLaMA-style decoders (1,600+ languages per the public release).
@@
Open-weights models ran on a single NVIDIA H100 80GB SXM via RunPod ($\sim$6 GPU hours, $\sim$\$25). Azure ran via API ($\sim$\$5).
@@ 4 Results
Table 1: Common Voice ga-IE (874 utterances). ABAIR is self-reported (Lonergan et al., 2025) using a different normaliser. Whisper results are zero-shot.
@@
ref ABAIR / Fotheidil closed w2v2-BERT 19.6*
@@
2 omniASR LLM 7B open w2v2+LLaMA 30.6 25.0 2.5 3.2 14.6
@@
4 omniASR LLM 300M open w2v2+LLaMA 37.6 29.3 3.4 4.9 19.2
@@
ref ABAIR / Fotheidil closed w2v2-BERT 44.5*
1 omniASR LLM 7B open w2v2+LLaMA 39.1 32.2 3.4 3.5 18.6
2 omniASR LLM 300M open w2v2+LLaMA 47.7 38.4 4.9 4.4 24.1
@@ 5.1 The Whisper variants we evaluated fail catastrophically on Irish
All four Whisper variants exceed 100% WER on both datasets, with insertion rates of 20–491%: the decoder emits fluent English unrelated to the input (Appendix C). v3 is worse than v2 (125.6% vs 106.0% on CV), and turbo is worse again (587.6% on FLEURS). The v2-large baseline matches the 110.4% WER reported by Qian et al. (2024); we did not find prior documentation of this monotonic v2 $\to$ v3 $\to$ turbo regression on Irish.
@@ 5.2 Common Voice overestimates performance
Models fine-tuned on Common Voice degrade 34–43 WER points on FLEURS (Table 3). Azure degrades 35 points. Models with massively multilingual pre-training degrade only 7–10 points. Prior work has evaluated across corpora (Lonergan et al., 2025), but the magnitude of this gap has not been quantified under a single harness. Common Voice WER alone is not a reliable proxy for deployment; BlasBench therefore requires evaluation on at least two datasets.
@@ 5.3 The open–closed gap
ABAIR’s Fotheidil reports its best result (10.9% WER) on private MíleGlór data, alongside 19.6% on Common Voice and 44.5% on FLEURS-R (Lonergan et al., 2025; Ma et al., 2024). [^4] ABAIR evaluates on FLEURS-R (re-recorded audio), not the original FLEURS we use, and with a different normaliser. Numbers are indicative, not directly comparable. omniASR 7B reaches 30.6% and 39.1% on original FLEURS under our normaliser. Because datasets and normalisers differ, we treat these numbers as indicative rather than a direct open-vs-closed measurement; they are consistent with a narrower gap than the headline 10.9%/30.6% suggests, but do not establish one. ABAIR’s architecture (wav2vec2-BERT 2.0) belongs to the same family as the strongest open models here, and their reported gains coincide with substantial Irish-specific data (398h labelled, 3,230h pseudo-labelled); isolating data vs. architecture as the cause is left to future work (Lonergan et al., 2025).
@@ 6 Related Work
Akhtar et al. (2024) benchmark Whisper, MMS, and SeamlessM4T on Urdu under a global-aggregation harness; BlasBench adopts the same methodology and adds bootstrap CIs and cross-corpus analysis. Klejch et al. (2025) build Scottish Gaelic ASR with Gaelic-specific normalisation, Jones (2022) fine-tune wav2vec2 for Welsh, and Bartley and Ragni (2025) demonstrate minimal-data ASR for Manx and Cornish from a spoken dictionary. On evaluation methodology, Manohar and Pillai (2024) show that multilingual ASR normalisers can distort error rates when they erase language-specific orthographic information. On statistical methodology, Gillick and Cox (1989) and Bisani and Ney (2004) remain the canonical references for significance testing of WER comparisons.
@@
Beyond MMS and Whisper, Zhang et al. (2023) (Google USM), Puvvada et al. (2024) (NVIDIA Canary), and Peng et al. (2024) (OWSM-CTC) all target broad multilingual speech coverage, but none of these papers establishes end-user Irish ASR performance. Multilingual benchmarks—FLEURS (Conneau et al., 2022a), ML-SUPERB (Shi et al., 2023), XTREME-S (Conneau et al., 2022b), and the Open ASR Leaderboard (Srivastav et al., 2025)—include Irish among many languages but apply no Irish-specific normalisation. Barry et al. (2022) release gaBERT for Irish NLP but not for ASR evaluation.
@@ 7 Conclusion
We evaluate on fixed public BlasBench evaluation slices for Common Voice ga-IE and FLEURS ga-IE because no larger public Irish test set is available to us; both are read speech, and no public conversational Irish ASR benchmark is available. The quoted counts (874 and 842) refer to those released BlasBench slices, not necessarily to every upstream packaging of these datasets. Neither dataset spans all three dialects equally or provides dialect labels—a gap in the Irish data landscape rather than a design choice. ABAIR’s numbers are self-reported on FLEURS-R (not original FLEURS) with a different normaliser, so the open–closed comparison is indicative, not exact.

### 16 Apr 2026 v2

@@ Abstract
Existing multilingual benchmarks include Irish among dozens of languages but apply no Irish-aware text normalisation, leaving reliable and reproducible ASR comparison impossible. We introduce BlasBench, an open evaluation harness that provides a standalone Irish-aware normaliser preserving fadas, lenition, and eclipsis; a reproducible scoring harness and per-utterance predictions released for all evaluated runs. We pilot this by benchmarking 12 systems across four architecture families on Common Voice ga-IE and FLEURS ga-IE. All Whisper variants exceed 100% WER through insertion-driven hallucination. Microsoft Azure reaches 22.2% WER on Common Voice and 57.5% on FLEURS; the best open model, Omnilingual ASR 7B, reaches 30.65% and 39.09% respectively. Models fine-tuned on Common Voice degrade 33–43 points moving to FLEURS, while massively multilingual models degrade only 7–10—a generalisation gap that single-dataset evaluation misses.
@@
Jyoutir Rajthanks: Corresponding author. Independent Researcher jyoutirraj@gmail.com          John Conway Independent Researcher john@blasapp.com
@@ 1 Introduction
State-of-the-art multilingual systems fail on out-of-distribution audio in Irish. Microsoft Azure reaches 22.2% WER on read speech (Common Voice ga-IE) and 57.5% WER on the more naturalistic FLEURS ga-IE. Every Whisper variant we evaluate (Radford et al., 2023) produces more tokens than the reference contains, exceeding 100% WER on both datasets (Tables 12). Irish presents a genuine difficulty: initial consonant mutations, eclipsis, and a spelling system that encodes grammatical contrasts mean that language-agnostic normalisation silently discards linguistically meaningful variation.
The problem is not that nobody has worked on Irish ASR. The ABAIR group at Trinity College Dublin has produced a sustained research programme culminating in their Fotheidil system at 10.9% WER (Lonergan et al., 2025). Qian et al. (2024) fine-tune Whisper on Irish. Community wav2vec2 fine-tunes exist on HuggingFace (Faste, 2022). The IWSLT shared task has produced Irish–English speech translation systems (Ahmad et al., 2024; Moslem, 2024). Multilingual benchmarks—FLEURS (Conneau et al., 2022a), MMS (Pratap et al., 2023), ML-SUPERB (Shi et al., 2023)—include Irish among dozens or hundreds of languages. These efforts span architecture families and data regimes, yet they have not been placed on a common footing.
The gap is methodological rather than empirical. ABAIR’s best result is on private MíleGlór data; their published Common Voice and FLEURS-R figures cover only their own systems. Community models self-report on different Common Voice versions with different normalisers. Multilingual benchmarks include Irish without Irish-specific normalisation or targeted analysis. [^1] ML-SUPERB and XTREME-S (Conneau et al., 2022b) include Irish data points but do not compare end-user ASR systems on Irish specifically. We did not identify a prior open, Irish-specific comparison of end-user ASR systems across architecture families under a shared Irish-aware evaluation protocol.
We present BlasBench, [^2] Harness: github.com/jyoutir/blasbench. Per-utterance predictions and aggregate results for every run in this paper are released as v0.1.0. an open benchmark and evaluation harness for Irish ASR. The harness applies Irish-aware text normalisation before scoring, so that orthographic variants reflecting mutation and dialect do not inflate error rates artificially. A 12-model comparison across commercial, large-scale multilingual, and fine-tuned systems constitutes the first use of the benchmark; reproducible, language-aware scoring is also a prerequisite for any automated research pipeline that requires a stable signal. Dedicated low-resource benchmarks for Urdu (Arif et al., 2025) and Scottish Gaelic (Klejch et al., 2025) have made the same case for their respective languages; BlasBench brings it to Irish.
@@ 2.3 Evaluation pipeline
    Scoring. WER and CER are computed via global aggregation: total substitutions, insertions, and deletions across all utterances divided by total reference units. This corpus-level approach avoids the bias of per-sentence averaging and matches the convention used in low-resource benchmarking work (Arif et al., 2025).
@@
    Confidence intervals. Bootstrap 95% CIs (Bisani and Ney, 2004) are computed with 1,000 resamples at a fixed seed (42), resampling utterance-level error counts and recomputing the global aggregate. CIs for every reported point estimate are released alongside the predictions in each run’s results.json; tables show point estimates for compactness.
@@ 2.4 Irish-aware normalisation
Standard ASR normalisers lowercase text, strip punctuation, and collapse whitespace. Recent work on multilingual ASR evaluation shows that language-agnostic normalisation can materially distort reported error rates (Manohar et al., 2024). For Irish, it destroys linguistically significant information.
@@
ABAIR’s pipeline includes text-processing components used inside Fotheidil (Lonergan et al., 2025) but no Irish ASR normaliser has been released standalone. To our knowledge, BlasBench ships the first standalone open-source Irish ASR evaluation normaliser released specifically for Irish ASR scoring; unreleased internal tools may exist.
@@ 3 Systems Evaluated
Whisper (Radford et al., 2023): Encoder-decoder transformer. Irish is absent from Whisper’s language list across v1, v2, and v3,333ga is absent from the LANGUAGES dictionary in whisper/tokenizer.py for v1/v2/v3, and the v3 model card provides no per-language training-data disclosure that would indicate Irish coverage. so results here are effectively zero-shot. We test medium (769M), large-v2 (1.5B), large-v3 (1.5B), and large-v3-turbo (809M).
@@
Meta multilingual: MMS-1B-All (Pratap et al., 2023) (1,107 languages); Omnilingual ASR (omniASR) LLM (Keren et al., 2025), wav2vec2 encoders with LLM-inspired Transformer decoders covering 1,600+ languages, released in 300M, 1B, 3B, and 7B sizes; we test the 300M and 7B endpoints.
@@
Open-weights models ran on a single NVIDIA H100 80GB SXM via RunPod ($\sim$6 GPU hours, $\sim$\$25). Azure ran via API ($\sim$\$5 at April 2026 pricing). Whisper, wav2vec2, MMS, and Azure runs all used the released harness; the omniASR runs used a separate fairseq2 driver at the time of writing, with a harness adapter pending.
@@ 4 Results
Table 1: Common Voice ga-IE (874 utterances). ABAIR is self-reported (Lonergan et al., 2025) using a different normaliser. Whisper results are zero-shot. Type column: open = weights and training data public; open-w = weights public, training data not fully disclosed; API = commercial cloud API; closed = weights unavailable.
@@
ref ABAIR / Fotheidil closed TDNN-HMM 19.6*
@@
2 omniASR LLM 7B open-w w2v2+Trans 30.6 25.0 2.5 3.2 14.6
@@
4 omniASR LLM 300M open-w w2v2+Trans 37.6 29.3 3.4 4.9 19.2
@@
ref ABAIR / Fotheidil closed TDNN-HMM 44.5*
1 omniASR LLM 7B open-w w2v2+Trans 39.1 32.2 3.4 3.5 18.6
2 omniASR LLM 300M open-w w2v2+Trans 47.7 38.4 4.9 4.4 24.1
@@ 5.1 The Whisper variants we evaluated fail catastrophically on Irish
All four Whisper variants exceed 100% WER on both datasets, with insertion rates of 20–491%: the decoder emits fluent English unrelated to the input (Appendix C). v3 is worse than v2 (125.6% vs 106.0% on CV), and turbo is worse again (587.6% on FLEURS). Qian et al. (2024) report 110.4% WER for whisper-large-v3 on FLEURS Irish, which is in the same regime as our results; we did not find prior documentation of the monotonic v2 $\to$ v3 $\to$ turbo regression on Irish.
@@ 5.2 Common Voice overestimates performance
Models fine-tuned on Common Voice degrade 33–43 WER points on FLEURS (Table 3). Azure degrades 35 points. Models with massively multilingual pre-training degrade only 7–10 points. Prior work has evaluated across corpora (Lonergan et al., 2025), but the magnitude of this gap has not been quantified under a single harness. Common Voice WER alone is not a reliable proxy for deployment; BlasBench therefore requires evaluation on at least two datasets.
@@ 5.3 The open–closed gap
ABAIR’s Fotheidil reports its best results (10.9% WER on private MíleGlór, 19.6% on Common Voice, and 44.5% on FLEURS-R) using their full pipeline including RNNLM rescoring; without LM rescoring the M1 acoustic model alone reports 23.7% on Common Voice (Lonergan et al., 2025; Ma et al., 2024). [^4] ABAIR evaluates on FLEURS-R (audio restored from the original FLEURS via the Miipher model), not the original FLEURS we use, and with a different normaliser. Numbers are indicative, not directly comparable. omniASR 7B reaches 30.6% and 39.1% on original FLEURS under our normaliser. Because datasets and normalisers differ, we treat these numbers as indicative rather than a direct open-vs-closed measurement; they are consistent with a narrower gap than the headline 10.9%/30.6% suggests, but do not establish one. ABAIR’s architecture (TDNN-HMM, Kaldi-based) is from a different family than the strongest open models here (wav2vec2-based), and their reported gains coincide with substantial Irish-specific training data (398h labelled, 3,230h pseudo-labelled (Lonergan et al., 2025)) that exceeds what any open model was trained on for Irish. The primary binding constraint on Irish ASR is data, not architecture.
@@ 6 Related Work
Arif et al. (2025) benchmark Whisper, MMS, and SeamlessM4T on Urdu with corpus-level WER aggregation; BlasBench adopts a similar aggregation strategy and adds bootstrap CIs and cross-corpus analysis. Klejch et al. (2025) build Scottish Gaelic ASR with Gaelic-specific normalisation, Jones (2022) fine-tune wav2vec2 for Welsh, and Bartley and Ragni (2025) demonstrate minimal-data ASR for Manx and Cornish from a spoken dictionary. On evaluation methodology, Manohar et al. (2024) show that multilingual ASR normalisers can distort error rates when they erase language-specific orthographic information. On statistical methodology, Gillick and Cox (1989) and Bisani and Ney (2004) remain the canonical references for significance testing of WER comparisons.
@@
Beyond MMS and Whisper, Zhang et al. (2023) (Google USM), Puvvada et al. (2024) (NVIDIA Canary), and Peng et al. (2024) (OWSM-CTC) all target broad multilingual speech coverage, but none of these papers establishes end-user Irish ASR performance. Multilingual benchmarks FLEURS (Conneau et al., 2022a), ML-SUPERB (Shi et al., 2023), and XTREME-S (Conneau et al., 2022b) include Irish among many languages but apply no Irish-specific normalisation; the Open ASR Leaderboard (Srivastav et al., 2025) does not include Irish at all (its multilingual track currently covers German, French, Italian, Spanish, and Portuguese). Barry et al. (2022) release gaBERT for Irish NLP but not for ASR evaluation.
@@ 7 Conclusion
We evaluate on fixed public BlasBench slices for Common Voice ga-IE (874 utterances) and FLEURS ga-IE (842 utterances); both consist of read speech, and no public conversational Irish ASR benchmark exists. These sample sizes limit the precision of pairwise comparisons. Confidence intervals are reported in the accompanying JSON; main tables show point estimates for compactness, as noted in the evaluation pipeline.
Neither dataset spans all three dialects equally, and neither provides dialect labels, reflecting a gap in available Irish data rather than a design choice. Dialect-stratified evaluation is not possible with current public resources.
The 12 systems evaluated cover the main open, community, and commercial categories; selection was constrained by licence, cost, and available credentials. OmniASR was run outside the shared evaluation harness because a fairseq2 driver dependency was unresolved at the time of writing, with a harness adapter pending; its numbers are therefore not directly comparable at the implementation level. Azure results depend on the API version and model snapshot available at time of evaluation (April 2026); future versions may differ.
ABAIR uses FLEURS-R rather than original FLEURS as its test set, and its normaliser is not publicly available; the open–closed comparison is therefore indicative rather than controlled. The WER impact of our normalisation choices has not been quantified.
The community wav2vec2 models were fine-tuned on earlier Common Voice releases (CV 6.1–8.0) and evaluated on CV 25.0. Some of the cross-corpus gap likely reflects distribution shift alongside genuine generalisation behaviour, with the multilingual-pretrained models in Table 3 providing a partial control.